OpenStudy (studio-games):

Given the function f(x) = 2(3)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other.

OpenStudy (studio-games):

Ill medal and fan you if you help out plz.

OpenStudy (solomonzelman):

Your function is: \(\color{blue}{f(x)=2(3)^x}\) (just verifying) \(\left\{\rm Step~~\text{#}1\right\}\) Find the following quantities for me please. f(0), f(1), f(2), f(3). \(\left\{\rm Step~~\text{#}2\right\}\) Use the slope formula, \({\rm m}=(y_2 -y_1)/(x_2-x_1)\) (for each section).

OpenStudy (studio-games):

Look I am horrible at math.

OpenStudy (studio-games):

So give me a minute please.

OpenStudy (solomonzelman):

sure, take your time...

OpenStudy (studio-games):

Oh ok I understand it now! Thanks! Gtg now my father needs me to do some chores.

OpenStudy (studio-games):

Medalled and fanned you.

OpenStudy (solomonzelman):

Not a must-do, but I certainly don't mind \(\Huge ☻\) yw